# A Longitudinal, Multinational, and Multilingual Corpus of News Coverage of the Russo-Ukrainian War

### v1

@@ 4 Data Annotations
4.1 Text Processing for Annotation
@@
Publication dates are parsed and normalized to ISO 8601 standard (yyyy-mm-dd), and multiple metadata fields—e.g., “publication date”, “published on”—are mapped to metadata fields in our unified schema (see Table 2). Text content is stripped of non-informative boilerplate elements like location or date prefixes commonly found in news articles (e.g., “KYIV, March 3 –”), inline footnotes, bylines, and repeated source disclaimers. Duplicate articles are removed, to avoid redundancy in print syndication and multilingual reprints. To support downstream linguistic analyses, the text is then converted to UTF-8 encoding, segmented into paragraphs, and split into sentences.
@@
economy Macroeconomic issues and markets: sanctions/trade, inflation, growth, jobs, taxes, energy/commodity prices.
@@
identity Group identity and belonging: nationality, ethnicity, religion, language, or gender; discrimination or identity politics.
@@
morality Explicit ethical/value framing: moral duty, fairness, dignity, humanitarian arguments.
@@
legal Courts and legal processes: rulings, lawsuits, prosecutions, designations, warrants, legal rights or entitlements.
@@
policy Concrete government outlook or outreach, rules, bills, decrees, program design and details (what a policy does).
@@
politics elections, party/elite competition, horse-race strategy, reshuffles, coalition dynamics.
@@
public opinion Polls, surveys, attitudes and protests as expressions of public sentiment.
@@
security War/defense, violence/crime, policing, border enforcement, terrorism, cyberattacks.
@@
welfare Benefits and human services: healthcare, pensions, housing aid, unemployment, assistance to refugees or internally displaced people.
Table 3: Topical frame definitions in dnipro: a significant abridgment of the Refugees and Migration Framing Vocabulary Yu and Fliethmann (2022).
4.2 Machine Translation
@@
A quarter of dnipro consists of Mandarin Chinese and Russian texts, (zh and ru, respectively, in ISO 639-1). We translate these into English (en) to facilitate downstream analyses. We choose three models based on recent comparative studies on neural machine translation by Smirnov et al. (2022) and Alemayehu et al. (2024): OPUS-MT Tiedemann et al. (2023), M2M100 Fan et al. (2020), and NLLB-200-Distilled Team (2022). Since there were no readily available explicit benchmarks for zh-to-en or ru-to-en translations, we select a random set of ten articles from Xinhua, evaluated by two native speakers of Mandarin Chinese for semantic fidelity and fluency. The evaluators worked independently to score the output of these models on 1-5 Likert scale for fidelity and fluency, and M2M100-418M received the highest scores from both evaluators on both criteria. So, we employ this model for all zh-to-en translations. Russian articles were similarly translated using the ru-to-en path of M2M100-418M.
4.3 Annotation Methodology
@@
To enable fine-grained analysis of key actors and locations, we perform Named Entity Recognition (NER) across the entire corpus (after translation to English). This annotation is carried out using the spaCy library Honnibal et al. (2020) and its pretrained model en_core_web_sm, identifying entities and their types according to the OntoNotes 5.0 taxonomy Weischedel et al. (2013). This process results in the identification of 85,155 distinct named entity mentions, and provides a foundational layer for tasks such as tracking the prominence of specific actors over time, temporal studies of media focus across entities and events, and entity-specific sentiment and stance analyses.
@@
To capture the sentiment of some of the most prominent named entities, we provide sentence-level sentiment annotation using a pretrained model: sentiment-roberta-large-english Hartmann et al. (2023). For each article, sentences containing such entities (e.g., Kursk, Bucha, Kherson, Zelensky, Ukraine, Russia) are identified using spaCy, and passed to this classifier. The model outputs are then mapped to continuous sentiment scores, which are then aggregated on a daily and monthly basis at both the source and country level using source–country relations. These annotations thus provide time series data revealing sentiment trends for each entity.
The dataset includes stance annotations obtained by few-shot instruction prompting with Phi-4 14B model  Microsoft (2024), identifying the stance towards a specified target. The input comprises a sentence and the immediate context in its news article. As part of the input, we provide a task instruction with a few examples, along with (i) a passage for context resolution, (ii) a focal sentence, and (iii) the target. The model is prompted to output a compact, formatted response:
@@
For topical framing, we again use few-shot prompting with Phi-4 14B model, where the input is an instruction describing the task, along with label definitions and 1-2 examples per label. Then, for a given instance – an article’s title, subtitle (if present) and content – we prompt the model to generate a json object containing the list of applicable topical frames, and associated confidence scores.
4.4 Human Evaluation
To evaluate the reliability of our automated annotations, we conduct human evaluations of the stance and topical frame annotations.
Frame PABAK Confidence
Human/Human Human/LLM Human LLM
Economy 0.70 (84.8%) 0.736 0.84 0.93
Identity 0.74 (86.8%) 0.783 0.79 0.92
Legal 0.85 (92.4%) 0.783 0.74 0.91
Morality 0.75 (87.3%) 0.750 0.79 0.91
Policy 0.55 (77.7%) 0.526 0.78 0.89
Politics 0.48 (74.1%) 0.425 0.79 0.92
Public Opinion 0.87 (93.4%) 0.865 0.80 0.89
Security 0.69 (84.3%) 0.587 0.85 0.93
Welfare 0.81 (90.4%) 0.783 0.80 0.87
Average 0.71 (85.7%) 0.693 0.798 0.909
Table 5: Inter-annotator agreement and confidence by topical frame. Pabak columns report human inter-annotator agreement (percent agreement in parentheses) and human–model agreement. Confidence columns report average confidence scores from human annotators and LLM (Phi-4).
Topical Frames:
Two expert annotators independently label a stratified random sample of 200 articles, providing nine binary labels with corresponding [0,1] confidence scores for each article. Significant class imbalance is a common challenge in multi-label annotation, and the inter-annotator agreement must consider the high base rate agreement where annotators agree on a label being absent from an article. The average agreement on label absence, over all nine topical frames, accounts for 65.82% of our evaluation sample, with the positive class (for any label) representing a minority of instances. Due to this prevalence bias, traditional measures like Cohen’s or Fleiss’ kappa are overly pessimistic reliability measures for our data. For a more accurate assessment that accounts for this imbalance, we report the prevalence-adjusted bias-adjusted kappapabak Byrt et al. (1993) – which is especially crucial for evaluation of consensus on the rarer labels.
The inter-annotator agreement results, measured by pabak, are presented in Table 5 and demonstrate substantial agreement on average (0.71, with 85.7% agreement). Three topical frame categories — Welfare, Public Opinion, and Legal — reveal near-perfect agreement (>> 0.8). Overall, seven of the nine labels achieve pabak scores $\geq$ 0.7, indicating a consistent and shared understanding of the annotation framework. The two remaining classes, Policy and Politics, have moderate agreement, as they represent more abstract and often overlapping conceptual domains. We also examine the correlation between the average confidence scores reported by the human experts and those reported by Phi-4. The framing categories are complex and sometimes overlapping, resulting in a weak positive correlation $\rho$ = 0.37.
Implications for dataset use. The strong agreement validates the overall reliability of our annotation schema. However, we recommend that users consider the tiered nature of this reliability when designing their experiments:
  • The seven frames with pabak > 0.69 are suitable for fine-grained analysis, including temporal tracking and cross-source comparison.
  • The Policy and Politics labels yield lower agreement due to their inherent conceptual overlap. We recommend these labels be used for high-level topic modeling or as supplementary features in larger models, with the understanding that they represent broader and possibly intersecting, categories of discourse.
  • The human-annotated sample of 200 documents, while modest in size, provides a high-quality validated benchmark. It was strategically selected through stratified sampling to ensure longitudinal representation across sources and languages, making it suitable for evaluating the performance of topical frame classification systems on real-world data on contentious wartime narratives.
@@ 4.4 Human Evaluation
Table 6: Inter-annotator agreements for stance labels, measured by percentage of agreement, pairwise Cohen’s $\kappa$, and Fleiss’ $\kappa$.
Stance Annotations:
@@
The stance annotation task identifies whether a sentence (with context) expressed a stance (favor or against) towards any entity, or if no stance was present (none). Two human annotators with native proficiency in English annotated a sample of 150 instances. We evaluate their inter-annotator agreement, and then also with phi-4 as the LLM annotator. Table 6 presents the pairwise and overall inter-annotator agreements, which indicate strong human inter-annotator agreement ($\kappa$ = 0.77). The LLM annotator achieved moderate agreement with both ($\kappa$ = 0.62 and 0.58), showing consistent but less-than-human reliability. Fleiss’ $\kappa$ = 0.643, however, suggests that the stance annotations are sufficiently reliable for large-scale analysis, despite the gap between human and automated annotation quality for this task.
Figure 2: Topical framing by source, with % share of articles for each label (confidence >> 0.5). Source abbr.: Financial Times (FT), Kyiv Post (KP), Sputnik (SP), European Pravda (EP), Global Times (GT), China Daily (CD), Xinhua (XH). Labels: (Sec)urity, (Pol)itics, (Econ)omy, (Welf)are, (Ident)ity, (Moral)ity, and (Pub)lic (Op)inion.
5 Use Case Experiments
@@
To demonstrate the analytical potential of dnipro, we present several studies that leverage its unique multi-perspective structure. These experiments are intended as use case demonstrations of the dataset’s capacity to support diverse research questions in computational journalism, political science, and discourse analysis. Each application highlights specific affordances of the corpus while suggesting broader research directions.
Thematic Analysis of Conflict Events:
@@
Our analyses focus on four major conflict events that received substantial global coverage and generated clear geopolitical divisions, inviting transnational comparisons (see Table 4). For each event, we filter articles using specific location identifiers (e.g., Bucha) to ensure event-specific analysis.
@@
To examine how media outlets prioritize different themes when covering conflict events, we analyze the topical content and emphasis patterns across sources. We identify which concepts, actors, and subtopics receive attention in coverage of major events, revealing how outlets differentially emphasize aspects such as historical context, humanitarian concerns, or institutional responses.
Using GPT-3.5 Ye et al. (2023) with structured prompts, we tag articles from weekly time windows surrounding each major event to identify key concepts, actors, and subtopics. For the events described in Table 4, we obtain 810 articles, with 10.5 such tags (avg.) per article. We aggregate these tags by event and source country to find the distinct thematic emphasis patterns across media outlets.
@@
Russian and Chinese outlets consistently emphasize institutional terminology (special military operation), security concerns (nuclear weapons), and counter-narratives (media manipulation, ceasefire), reflecting strategic adversarial framing. Ukrainian sources predominantly highlight humanitarian dimensions (civilian casualties) with frequent references to specific conflict locations. American and British media balance humanitarian topics with institutional responses (sanctions, international criminal court) and policy discussions (no fly zone).
@@
These patterns demonstrate how dnipro captures differences in thematic prioritization across geopolitical actors even when covering the same underlying events.
Russian Media Ukrainian Media
Russia had testimony from residents of Bucha and Mariupol, who reported brutal crimes, such as torture, committed by Ukrainian nationalists. (TASS) All eyewitnesses noted the systematic nature of the killings in Bucha by the Russian army. (Censor)
Russian forces jointly with border guards repelled these attacks and hit enemy positions near the border and in the Sumy Region. (TASS) Russian forces attacked six border areas and settlements of Sumy Oblast on June 7, firing 14 times and causing at least 54 explosions, the Sumy Oblast Military Administration reported. (Kyiv Independent)
Table 7: Contradictory claims from opposing media outlets, highlighting how conflicting narratives are constructed not through direct refutation, but via divergent attribution, framing, and selection biases.
Topical Framing:
@@
We conduct a corpus-wide analysis using our frame definitions (Table 3). Article-level topical frames provide a comparable map of what each media outlet emphasizes, revealing systematic priorities that contextualize downstream analyses such as stance, sentiment, or discourse. Articles are annotated with phi-4 using a few-shot instruction prompt, producing multi-label assignments along with confidence scores.
@@
Figure 2 shows that Security dominates across all sources, with the highest shares in Ukrainian media, followed by Russian state media and Western sources like CNN. The second-most prevalent category is Politics, elevated in Russian and Chinese sources. It is substantially present in other sources too. Economy is most visible in Chinese media and, naturally, The Financial Times, while Policy appears steadily across sources, with higher emphasis in Chinese and British coverage. Relative to others, Russian media leans more on Legal topics. Notably, discussions of Morality are extremely rare despite the focal topic of this dataset being an extended war, indicating that operational, strategic, and material considerations far outweigh explicit ethical rhetoric in war coverage.
Figure 3: Daily average of stance scores towards Ukraine, grouped by media sources from five nations. The first 15 days of the war (left), contrasted with the last 15 days collected for dnipro (right).
Stance analysis:
Based on our stance annotations in § 4.3, we provide a comparison of the stance towards Ukraine between the first fortnight of the war and the last 15 days of dnipro data. This reveals a striking temporal shift across geopolitical lines. In Figure 3, we see that Chinese media was slightly pro-Ukrainian. By August 2024, however, this shifted to mild opposition. Russian coverage is consistently negative, but there is a drastic negative shift. The US media stance becomes even more supportive, while UK news begins to exhibit greater variability. This temporal comparison demonstrates how dnipro can provide analytical insights into reporting patterns. In this use case, it illustrates the polarization of geopolitical alignments.
Divergent Narratives:

### v2

@@ 4.  Data Annotations
4.1.  Text Processing for Annotation
@@
Publication dates are parsed and normalized to ISO 8601 standard (yyyy-mm-dd), and multiple metadata fields—e.g., “publication date”, “published on”—are mapped to metadata fields in our unified schema (as shown in Table 2). The article text content is stripped of boilerplate elements like location or date prefixes commonly found in news articles (e.g., “KYIV, March 3 –”), inline footnotes, bylines, and repeated source disclaimers. Duplicate articles are removed, to avoid redundancy in print syndication and multilingual reprints. To support downstream linguistic analyses, the text is then converted to UTF-8 encoding, segmented into paragraphs, and split into sentences.
@@
economy 1 Macroeconomic issues and markets: sanctions/trade, inflation, growth, jobs, taxes, energy/commodity prices.
@@
identity 3 Group identity and belonging: nationality, ethnicity, religion, language, or gender; discrimination or identity politics.
@@
morality 2 Explicit ethical/value framing: moral duty, fairness, dignity, humanitarian arguments.
@@
legal 1 Courts and legal processes: rulings, lawsuits, prosecutions, designations, warrants, legal rights or entitlements.
@@
policy 1 Concrete government outlook or outreach, rules, bills, decrees, program design and details (what a policy does).
@@
politics 1 Political proceedings and party competition: elections, horse-race strategy, reshuffles, coalition dynamics.
@@
public opinion 1 Public attitudes and moods: polls, surveys, and protests as expressions of public sentiment and interest.
@@
security 1 War/defense, violence/crime, policing, border enforcement, terrorism, cyberattacks.
@@
welfare 1 Benefits and human services: healthcare, pensions, housing aid, unemployment, assistance to refugees or displaced people.
Table 3: Topical frames in dnipro: an abridgment of the Refugees and Migration Framing schema Yu and Fliethmann (2022), with its utilitarian 1, moral-universal 2, and identity-related 3 categories derived from the discourse ethics of Habermas (1991).
4.2.  Machine Translation
@@
A quarter of dnipro consists of Mandarin Chinese and Russian texts, (zh and ru, respectively, in ISO 639-1). We translate these into English (en) to facilitate downstream analyses. We choose three models based on recent comparative studies on neural machine translation by Smirnov et al. (2022) and Alemayehu et al. (2024): OPUS-MT Tiedemann et al. (2023), M2M100 Fan et al. (2021), and NLLB-200-Distilled Team (2022). Since there were no readily available explicit benchmarks for zh-to-en translation of war or conflict narratives, we select a random set of ten articles from Xinhua, evaluated by two native speakers of Mandarin Chinese for semantic fidelity and fluency. The evaluators worked independently to score the output of these models on 1-5 Likert scale for fidelity and fluency, and M2M100-418M received the highest scores from both evaluators on both criteria. So, we employ this model for all zh-to-en translations. Russian articles from Izvestia are similarly evaluated by native Russian speakers, and translated using the ru-to-en path of M2M100-418M.
4.3.  Annotation Methodology
@@
Named Entity Recognition: To enable fine-grained analysis of key actors and locations, we perform named entity recognition (NER) across the entire corpus (after translation to English). This annotation is carried out using spaCy Honnibal et al. (2020) and its pretrained model en_core_web_sm, identifying entities and their types according to the OntoNotes 5.0 taxonomy (ontonotes). This identifies 85,155 distinct named entity mentions, and provides the foundational layer for downstream tasks such as tracking the prominence of specific actors over time, temporal studies of media focus across entities and events, and entity-specific sentiment and stance analyses.
Sentiment Analysis: To capture the sentiment surrounding the most prominent named entities, we provide sentence-level sentiment annotation using a pretrained model: sentiment-roberta-large-english Hartmann et al. (2023). For each article, sentences containing such entities (e.g., Kursk, Bucha, Kherson, Zelensky, Ukraine, Russia) are identified using spaCy, and passed to this classifier. The model outputs are then mapped to continuous sentiment scores, which are then aggregated on a daily and monthly basis at both the source and country level using source–country relations. These annotations thus provide time series data revealing how sentiment toward specific actors and events evolves across geopolitical perspectives.
@@
Stance Detection: The dataset includes stance annotations obtained by few-shot instruction prompting with Phi-4 14B model  Microsoft (2024), identifying the stance towards a specified target. The input comprises a sentence and the immediate context in its news article. To Phi-4, we provide the task instruction with a few examples, along with (a) a passage for contextual information, (b) a focal sentence, and (c) the target. The model is prompted to output a compact, formatted response:
@@
Topical Framing: For topical framing, we use the same few-shot prompting with Phi-4 14B (Microsoft (2024)), where the input is an instruction describing the task, along with topical frame label definitions (shown in Table 3) and 1-2 examples per label. Then, for a given instance—comprising an article’s title, subtitle (if present) and content—we prompt the model to generate a json object containing the list of applicable topical frames and their corresponding confidence scores. This supports analysis of how different sources emphasize humanitarian, military, diplomatic, or economic dimensions of the conflict.
$4.$  Human Evaluation
To assess annotation quality, we conduct rigorous human evaluation on both stance detection and topical framing. Results demonstrate that automated labeling substantially agrees with expert judgment (pabak = 0.71 for topical framing, Fleiss’ $\kappa$ = 0.643 for stance detection), validating their reliability for coarse-grain analyses while identifying specific categories that warrant careful interpretation.
@@
Table 5: Annotation agreement on stance labels: agreement percentage, pairwise Cohen’s $\kappa$, and Fleiss’ $\kappa$.
4.4.1.  Stance Annotations
@@
The stance annotation task identifies whether a sentence (with context) expressed a stance (favor or against) towards any entity, or if no stance was present (none). Two human annotators with native proficiency in English annotated a sample of 150 instances. We evaluate their inter-annotator agreement, and then also with phi-4 as the LLM annotator. Results in Table 5 present the pairwise and overall inter-annotator agreements, demonstrating strong human inter-annotator agreement ($\kappa$ = 0.77) and moderate LLM-human agreements ($\kappa$ = 0.62 and 0.58), showing consistent but less-than-human reliability. The overall Fleiss’ $\kappa$ = 0.643 indicates that the stance annotations are substantially reliable for large-scale analysis.
While the LLM shows a consistent gap with human judgment, this is expected for nuanced pragmatic language tasks like stance detection. The annotations remain valuable for trend analysis and comparative studies across sources—such as identifying systematic differences in how Russian vs. Ukrainian sources frame political leadership or specific events—while nuanced high-stakes stance analyses may benefit from human review.
Frame pabak Confidence
Human All Human Phi-4
High reliability (pabak > 0.8)
Public Opinion 0.87 (93.4%) 0.865 0.80 0.89
Legal 0.85 (92.4%) 0.783 0.74 0.91
Welfare 0.81 (90.4%) 0.783 0.80 0.87
Substantial reliability (pabak 0.69 - 0.8)
Morality 0.75 (87.3%) 0.750 0.79 0.91
Identity 0.74 (86.8%) 0.783 0.79 0.92
Economy 0.70 (84.8%) 0.736 0.84 0.93
Security 0.69 (84.3%) 0.587 0.85 0.93
Moderate reliability (pabak < 0.69)
Policy 0.55 (77.7%) 0.526 0.78 0.89
Politics 0.48 (74.1%) 0.425 0.79 0.92
Average 0.71 (85.7%) 0.693 0.80 0.91
Table 6: Human inter-annotator agreement scores (pabak and percentage), human-model agreement (pabak), and annotator confidence scores ([0-1]) across all nine topical frames introduced in Table 3.
4.4.2.  Topical Frames
Two expert annotators independently label a stratified random sample of 200 articles, providing nine binary labels with corresponding [0,1] confidence scores for each article. Significant class imbalance is a common challenge in multi-label annotation, and the inter-annotator agreement must consider the high base rate agreement where annotators agree on a label being absent from an article. The average agreement on label absence, over all nine topical frames, accounts for 65.82% of our evaluation sample, with the positive class (for any label) representing a minority of instances. Due to this prevalence bias, traditional measures like Cohen’s or Fleiss’ $\kappa$ are overly pessimistic reliability measures for our data. For a more accurate assessment that accounts for this imbalance, we report the prevalence-adjusted bias-adjusted kappa, or pabak Byrt et al. (1993), which is especially crucial for evaluation of consensus on the rarer labels.
The inter-annotator agreement results, measured by pabak, are presented in Table 6 and demonstrate substantial agreement on average (0.71, with 85.7% agreement). Three topical frame categories — Welfare, Public Opinion, and Legal — reveal near-perfect agreement (>> 0.8). Overall, seven of the nine labels achieve pabak scores $\geq$ 0.7, indicating a consistent and shared understanding of the annotation framework. The two remaining classes, Policy and Politics, have moderate agreement, as they represent more abstract and often overlapping conceptual domains.
Further, we examine the correlation between the average confidence scores reported by the human experts and those reported by phi-4. The framing categories are complex and sometimes overlapping, resulting in a weak positive correlation $\rho$ = 0.37. Notably, phi-4 exhibits systematically higher confidence (avg. 0.91) than human annotators (avg. 0.80) across all categories, regardless of the extent of agreement. This suggests that users should not interpret phi-4’s self-reported confidence scores as reliable indicators of annotation quality.
Implications for Dataset Use: The strong agreement validates the overall reliability of our annotation schema. However, we recommend that users consider the following nuances in their studies:
  • 1)
    Frames with high/substantial reliability (pabak > 0.69) are suitable for fine-grained analysis such as tracking shifts in Welfare framing after major tragedies, or comparing Security or Legal discourse across state-aligned media.
  • 2)
    The moderately reliable frames (pabak < 0.69) — Policy and Politics — yield lower agreement due to their inherent conceptual overlap. These frame labels may be best suited for coarse-grained topic modeling or as supplementary features, with the understanding that they represent broader discourse categories that often intersect. As such, their unification may provide more reliable signals.
  • 3)
    The expert-annotated sample of 200 documents, stratified across sources, languages, and time periods, provides a completely validated high-quality gold-standard benchmark. Our stratified sampling ensures it is suitable for evaluating the performance of topical frame classification systems on real-world data on contentious wartime narratives.
Together, these evaluations demonstrate that dnipro’s annotations achieve a balance of scale and quality rare in wartime media corpora: automated methods enable comprehensive coverage, while stratified human validation ensures researchers can make informed decisions regarding labels, as per their task-specific goals. The annotation guidelines and task-specific prompts are provided in Appendix A and C, respectively.
5.  Use Case Experiments
@@
To demonstrate the analytical potential of dnipro, we present four use-case experiments exploiting its unique multi-perspective structure. These are not exhaustive analyses, but proof-of-concept demonstrations showing how the corpus enables diverse research in computational journalism, political communication, discourse analysis, and information warfare studies. Each application highlights specific affordances of the corpus while suggesting broader research directions.
Figure 2: Topical framing across news outlets (for each frame, % of articles with confidence > 0.5): Security dominates all outlets, while Economy is most prominent in Chinese media and Financial Times (ft). Other abbr.: Kyiv Independent (ki), Sputnik (sp), European Pravda (ep), Global Times (gt), China Daily (cd), Xinhua (xh).
@@
1. A Thematic Study of Conflict Events: Our analysis focuses on the four major conflict events shown in Table 4, which received substantial global coverage, generated clear geopolitical divisions, and invited transnational comparisons. For each event, we filter articles using specific location identifiers (e.g., Bucha) to ensure relevance.
@@
We examine how conflict events are framed by analyzing patterns of thematic salience across news outlets. By identifying which concepts, actors, and subtopics are underscored in coverage, we reveal how outlets differentially emphasize aspects such as historical context, humanitarian concerns, or institutional responses.
@@
Word clouds visualizing these thematic patterns for each conflict event and outlet country are provided in Appendix B. Our prelimnary analysis shows that Russian and Chinese outlets consistently emphasize institutional terminology (special military operation), security concerns (nuclear weapons), and counter-narratives (media manipulation, ceasefire), reflecting strategic adversarial framing. Ukrainian outlets predominantly highlight humanitarian dimensions (civilian casualties) with frequent references to specific conflict locations. Media from the USA and the UK balance humanitarian topics with institutional responses (sanctions, international criminal court) and policy discussions (no fly zone).
@@
Thus, dnipro captures differences in thematic salience and agenda-setting across geopolitical actors narrating the same underlying events.
@@
2. Topical Framing: We conduct a corpus-wide analysis using the schema introduced in Table 3. Article-level topical frames provide a comparable map of what each media outlet emphasizes, revealing systematic priorities that contextualize downstream analyses such as stance, sentiment, or discourse. Articles are annotated with phi-4 using a few-shot instruction prompt, producing multi-label assignments along with confidence scores (as described in § 4.4.2).
@@
Figure 2 shows that Security dominates across all outlets, with the highest shares in Ukrainian media, followed by Russian state media and Western outlets like CNN. The second-most prevalent category is Politics, elevated in Russian and Chinese media. It is substantially present in other outlets too. Economy is most visible in Chinese media and, naturally, The Financial Times, while Policy appears steadily across outlets, with higher emphasis in Chinese media and the UK. Relative to others, Russian media leans more on Legal topics. Notably, discussions of Morality are extremely rare despite the focal topic of this dataset being an extended war, indicating that operational, strategic, and material considerations far outweigh explicit ethical rhetoric in war coverage.
3. Stance Analysis: We compare stance toward Ukraine between the war’s first 15 days and the final 15 days of data collection (Figure 3), revealing striking temporal shifts. Chinese media shifted from mildly pro-Ukrainian to mildly oppositional, likely reflecting China’s deepening strategic alignment with Russia. Russian coverage moved from a mildly negative stance to overwhelmingly hostile – possibly correlating with battlefield changes and political rhetoric. The US coverage became even more supportive, while the UK exhibited greater variability, possibly due to the transition from early solidarity to war fatigue, domestic economic pressures and/or leadership changes.
The temporal dimension of dnipro proves crucial for these insights: China’s pivot, Russian and American intensification in opposite directions, and growing ambivalence in the UK.
Figure 3: Daily average stance scores toward Ukraine, grouped by nation, comparing the first 15 days of the war (left) with the last 15 days of data collection (right). Note the dramatic intensification of Russian negativity and American support, increased variability in the UK, and the gentle negative shift in Chinese media.